OpenStudy (anonymous):

PLEASE HELP: You borrowed $5000 from your parents to purchase a used car, You have agreed to make payements of $250 per month plus an additional 1% interest on the unpaid balance of the loan. a) is this an example of a geometric series or an arithmetic series? suppor answer mathematically. b)find the first year's monthly payments that you will make and the unpaid balance after each month. c) find the total amount of interest paid over the term of the loan.

OpenStudy (anonymous):

outstanding loan at the end of 1st month =5000-250=4750 1% means 47.5 additional payment 2nd month 4750-250=4500 1% means 45 3rd month 4500-250=4250 1% means 42.5 . so its an example of arithmetic progression

OpenStudy (anonymous):

you can calculate total yearly payment likewise